OpenStudy (anonymous):

What is the mid point. ? P(10, 2), Q(7, -3)

OpenStudy (anonymous):

x1+x2/2, y2+y1/2

OpenStudy (anonymous):

Okay i got \[\sqrt{137}\]

OpenStudy (anonymous):

is the right.?

OpenStudy (anonymous):

keep the two points separated so (x,y) is your answer since it's point. so it's not right

OpenStudy (anonymous):

So what would the distance formula for the would be.?

OpenStudy (anonymous):

it's midpoint not distance for your question. did you mean distance from start?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

for distance formula it's not correct

OpenStudy (anonymous):

im trying to find the mid point for P(10, 2), Q(7, -3)

OpenStudy (anonymous):

ok you where confusing, so you said yes to my question. so use the 1st formula y2-y1 then divde by 2 so 2+-3 = -1/2 = -.5 do same for x you get your point

OpenStudy (anonymous):

yes i was getting confuse lol

OpenStudy (anonymous):

you do same for x so x2-x1/2 gets you x point so (x,-.5)

OpenStudy (anonymous):

the x is 17/2 so midpoint is (17/2,-.5)